# DAC: Decomposed Automation Correction for Text-to-SQL

### v1

@@ 1 Introduction
Text-to-SQL is an important task, significantly reducing the overhead of obtaining information from the database by generating corresponding SQL based on user questions (Deng, Chen, and Zhang 2022). Specifically, about the text-to-SQL task, users provide the question and the related database, and then the model generates the corresponding SQL results. Currently, the methods based on Large Language Models (LLMs) have become the mainstream for text-to-SQL because LLMs can achieve higher performance than fine-tuned models, without fine-tuning (Hong et al. 2024b; Li et al. 2024a; Shi, Tang, and Yang 2024; Kanburoğlu and Tek 2024). Therefore, in this paper, we mainly focus on how to enhance the performance of the text-to-SQL task based on LLMs.
Recent research shows that automated correction is an effective method to improve text-to-SQL performance (Chen et al. 2024; Askari, Poelitz, and Tang 2024; Xie et al. 2024; Wang et al. 2024a), where LLMs first generate an initial SQL and then detect mistakes of the SQL as correction feedback (Pan et al. 2024). For instance, SQL-CRAFT (Xia et al. 2024) assesses whether the SQL execution results on databases satisfy the user question and corrects the SQL accordingly. EPI-SQL (Liu and Tan 2024) builds an example library of error cases and then retrieves similar cases from the library to guide each correction. However, previous studies indicate that it could be not effective for LLMs to directly correct all mistakes in the generated SQL since it is hard for LLMs to detect mistakes in the generated results (Zhang et al. 2024a; Kamoi et al. 2024; Zhang et al. 2024c).
@@
Previous works (Dhuliawala et al. 2024; Vacareanu et al. 2024) show that correction with the decomposed reasoning process has better performance than directly correcting because detecting and correcting the mistakes in the reasoning process is easier than in the final answer. Therefore, in this paper, we discuss that: (i) Experimentally, we present that correction with decomposed sub-tasks is more effective than direct correction; (ii) Methodologically, we propose correcting the text-to-SQL results based on the decomposed tasks of entity linking and skeleton parsing to enhance the performance of the automation correction.
First, we discuss that correction with the decomposed sub-tasks can bring better correction performance than the direct correction, as shown in Figure 1. We first present that decomposing into sub-tasks ensures better automation correction performance since mistakes can be more easily detected and corrected based on sub-tasks than using the original generation result. Applying the above conclusion, we propose decomposing text-to-SQL into two sub-tasks: entity linking and skeleton parsing, which are the two most crucial capabilities for addressing the text-to-SQL task (Li et al. 2023a; Deng, Chen, and Zhang 2022; Hong et al. 2024b).
Based on the above analysis, we propose Decomposed Automation Correction (DAC), which corrects the generated SQL with the mistakes of entity linking and skeleton parsing as feedback. Schema linking denotes detecting the table and column names related to the question, where we use LLMs to directly generate relevant entities based on the user question and the database, following Li et al. (2023c). Skeleton parsing denotes generating the SQL skeleton corresponding to the user question, where we employ LLMs to generate the SQL skeleton according to the prior research (Guo et al. 2024b, a). After generating the entities and skeleton, we take their inconsistencies with the initial SQL as feedback for the correction, as shown in Figure 1.
To validate DAC, we conduct experiments on three mainstream text-to-SQL datasets: Spider (Yu et al. 2018), Bird (Li et al. 2023b), and KaggleDBQA (Lee, Polozov, and Richardson 2021). The experiments show that our method achieves an average improvement of $3.7\%$ compared to baselines, and a $1.3\%$ improvement over previous text-to-SQL automation correction methods, demonstrating the effectiveness of DAC.
@@ 2.2 Adaption to Text-to-SQL
We apply the discussion above to the text-to-SQL task, where we use entity linking and skeleton parsing as sub-tasks for automated correction. Entity linking refers to identifying relevant table and column names from the given database schema based on the user question (Liu et al. 2021). Skeleton parsing denotes creating the corresponding SQL skeleton of the user question, which removes the table names, column names, and values in SQL (Guo et al. 2024b). We use these two sub-tasks because previous work regards them as the most crucial capabilities for text-to-SQL (Li et al. 2023a; Deng, Chen, and Zhang 2022; Hong et al. 2024b).
@@ 3.1 SQL Generation
Before the correction, we first generate an initial SQL to be corrected. Following previous work (Xia et al. 2024), we use few-shot learning to generate SQL. We use BM25 to select demonstrations similar to the user question from the given demonstration pool, together with the database and the user question as the input. With the input, LLMs directly generate a single SQL as the initial result to be corrected.
@@ 3.2 Entity Linking
After obtaining the initial SQL, we aim to identify mistakes in the linked entities. To achieve this, it is essential to know the entities related to the user question, specifically the table and column names in the question-related database. Following previous work (Andrew et al. 2024; Goel et al. 2023), we use LLMs to select the question-related tables and columns. We employ the tokenized user question and the database as inputs and output the entity linking information in the format of List[Dict[str, str]] in Python, where the length of the list matches the number of tokens in the tokenized question. We use this format because the performance of entity linking generated in programmatic format by LLMs is superior to only generating entities (Li et al. 2023c).
@@ 3.3 Skeleton Parsing
After identifying the linked entities, we seek to parse the skeleton corresponding to the user question to identify mistakes in the initial SQL generation on the skeleton. Previous research indicates that LLMs can parse the SQL skeleton based solely on the user question since the skeleton corresponds to the syntactic logic of the question without requiring database information (Gu et al. 2023; Pourreza et al. 2024; Kothyari et al. 2023). Therefore, we use only the user question as the input and ask LLMs to parse an SQL skeleton corresponding to the question. Subsequently, we remove entities from the parsed SQL, since these entities could not correspond to the column names or table names in the original database, thus obtaining the SQL skeleton.
@@ 4.1 Settings
Datasets
We use Spider (Yu et al. 2018), Bird (Li et al. 2023b), and KaggleDBQA (Lee, Polozov, and Richardson 2021) to evaluate DAC. Spider is one of the most mainstream text-to-SQL datasets, which covers multi-table and multi-domain questions. Compared with Spider, Bird is more difficult, and databases are more complex and more near to the practical application. The difficulty of KaggleDBQA is close to that of Bird, but the scale of the database is larger [^2] We call KaggleDBQA as Kaggle for simplicity..
Metric
In this paper, following the previous works (Askari, Poelitz, and Tang 2024; Cen et al. 2024), we use execution accuracy (EX) to evaluate our method. EX refers to the proportion of data that predicted SQL execution results are the same as correct SQL execution results.
Models
We verify our methods on the instruction fine-tuning version of Llama3 (Meta 2024), Deepseek-Coder (Deep.C.) (Guo et al. 2024c), and gpt-3.5-turbo-1106 (GPT-3.5) (Ye et al. 2023). Llama3 and Deepseek-Coder are the most mainstream general and code LLMs in the current open-source LLMs, respectively. While GPT-3.5 is one of the most popular closed-source LLMs. Therefore, our experimental LLMs can well cover different application scenarios.
Implementation Details
We employ few-shot inference for each step in DAC. Following the previous work (Chang and Fosler-Lussier 2023), we use 5-shot for the inference. We use the Spider training set as the candidate demonstration pool and use BM25 to select the most similar demonstrations for each user question. For the entity linking step, we employ the labeled linking information of Spider collected by Liu et al. (2021), which is used to guide the generation format (Madaan, Hermann, and Yazdanbakhsh 2023) and can be adapted to different databases as shown in Table 1. The prompts we used can be seen in the supplementary material.
@@ 4.2 Main Experiment
gpt-3.5-turbo Self-Debug (Chen et al. 2024) $71.1\rightarrow 72.2$-
SQL-CRAFT (Xia et al. 2024) $78.2\rightarrow 79.3$-
@@
gpt-4 Self-Debug (Chen et al. 2024) $73.2\rightarrow 73.6$-
EPI-SQL (Liu and Tan 2024) $79.7\rightarrow 85.1$-
SQL-CRAFT (Xia et al. 2024) $79.7\rightarrow 85.4$ $53.3\rightarrow 55.2$
MAGIC (Askari, Poelitz, and Tang 2024) $78.6\rightarrow\bm{85.7}$-
@@ 4.4 Analysis
Is SQL with the correct entity and skeleton the correct SQL?
@@
Can LLMs effectively generate entity linking information?
@@
To validate the performance of LLMs in entity linking, we conduct experiments on the schema linking of the Spider dev set. The experimental results are shown in Table 5, from which we can see that: (i) In most metrics, the performance of entity linking based on LLMs is superior to that based on grams or fine-tuning methods (Liu et al. 2021), demonstrating the effectiveness of LLMs in entity linking; (ii) Compared to code-specific LLMs, general LLMs exhibit better entity-linking performance in most metrics, indicating that general-purpose LLMs are more suitable for the entity-linking task due to the better instruction following.
Is the SQL skeleton generated not based on databases better than based on databases?
@@
What is the performance of DAC using oracle entities and skeletons?
@@
What is the performance of DAC on SQLs with different hardness?
@@ 5.1 Text-to-SQL
Research on the text-to-SQL task investigates how to generate corresponding SQL results based on given user questions and databases (Deng, Chen, and Zhang 2022). The past research is mainly based on small-scale models (Guo et al. 2019; Bogin, Gardner, and Berant 2019; Wang et al. 2020; Scholak, Schucher, and Bahdanau 2021; Dou et al. 2022) like BERT (Devlin et al. 2019) and BART (Lewis et al. 2020). Currently, the methods based on LLMs have become the mainstream of the text-to-SQL task since their brilliant performance without fine-tuning (Hong et al. 2024b). One type of approach is to fine-tune LLMs to help them better adapt to the text-to-SQL task by directly using the text-to-SQL data (Zhang et al. 2024b; Li et al. 2024b; Pourreza and Rafiei 2024), or merging the data of the tasks related to the text-to-SQL task (e.g., entity linking, knowledge generation) (Sun et al. 2024; Hong et al. 2024a). Another type of approach uses LLMs by few-shot inference, achieving significantly better performance without the overhead of fine-tuning, like in-context learning (Pourreza and Rafiei 2023; Gao et al. 2024; Chang and Fosler-Lussier 2023), self-consistency (Ni et al. 2023; Zhong et al. 2023; Lee et al. 2024), and multi-agent (Wang et al. 2024a; Xie et al. 2024).
@@ 5.2 Automated Correction
The automated correction approach first determines the correctness of the generated result after producing an answer, where if the answer is incorrect, it makes the necessary corrections (Pan et al. 2024). The most mainstream automated correction methods primarily focus on the code generation task, such as using feedback during training to correct errors (Ouyang et al. 2022; Feng et al. 2024; McAleese et al. 2024) or having the model judge and correct errors after inference (Chen et al. 2023, 2024; Wu et al. 2024; Zhang et al. 2024c). For the text-to-SQL task, automated correction methods pay more attention to the characteristics of the task itself, such as correcting based on execution results from the database (Li and Xie 2024; Zhong et al. 2023; Xia et al. 2024) or generating SQL-related error messages (Liu and Tan 2024; Askari, Poelitz, and Tang 2024).
However, previous work shows that LLMs struggle to directly identify SQL mistakes (Zhang et al. 2024a; Kamoi et al. 2024). Therefore, we propose DAC to provide decomposed feedback from the sub-tasks of entity linking and skeleton parsing, thereby helping correct mistakes.
@@ Appendix A Prompts of DAC
The prompts of DAC is shown in Table 9, Table 10, Table 11, Table 12, and Table 13. The format of “{schema}” in our prompts is the same as the definition in (Chang and Fosler-Lussier 2023).

### v2

@@ 1 Introduction
Text-to-SQL is an important task, significantly reducing the overhead of obtaining information from the database by generating corresponding SQL based on user questions (Deng et al., 2022). Specifically, about the text-to-SQL task, users provide the question and the related database, and then the model generates the corresponding SQL results. Currently, the methods based on Large Language Models (LLMs) have become the mainstream for text-to-SQL because LLMs can achieve higher performance than fine-tuned models, without fine-tuning (Hong et al., 2024b; Li et al., 2024a; Shi et al., 2024; Kanburoğlu & Tek, 2024). Therefore, in this paper, we mainly focus on how to enhance the performance of the text-to-SQL task based on LLMs.
Recent research shows that automated correction is an effective method to improve text-to-SQL performance (Chen et al., 2024; Askari et al., 2024; Xie et al., 2024; Wang et al., 2024a), where LLMs first generate an initial SQL and then detect mistakes of the SQL as correction feedback (Pan et al., 2024). For instance, SQL-CRAFT (Xia et al., 2024) assesses whether the SQL execution results on databases satisfy the user question and corrects the SQL accordingly. EPI-SQL Liu & Tan (2024) builds an example library of error cases and then retrieves similar cases from the library to guide each correction. However, previous studies indicate that it could be not effective for LLMs to directly correct all mistakes in the generated SQL since it is hard for LLMs to detect mistakes in the generated results (Zhang et al., 2024a; Kamoi et al., 2024; Zhang et al., 2024c).
@@
Previous works Dhuliawala et al. (2024); Vacareanu et al. (2024) show that correction with the decomposed reasoning process has better performance than directly correcting because detecting and correcting the mistakes in the reasoning process is easier than in the final answer. Therefore, in this paper, we discuss that: (i) Experimentally, we present that correction with decomposed sub-tasks is more effective than direct correction; (ii) Methodologically, we propose correcting the text-to-SQL results based on the decomposed tasks of entity linking and skeleton parsing to enhance the performance of the automation correction.
First, we discuss that correction with the decomposed sub-tasks can bring better correction performance than the direct correction, as shown in Figure 1. We first present that decomposing into sub-tasks ensures better automation correction performance since mistakes can be more easily detected and corrected based on sub-tasks than using the original generation result. Applying the above conclusion, we propose decomposing text-to-SQL into two sub-tasks: entity linking and skeleton parsing, which are the two most crucial capabilities for addressing the text-to-SQL task Li et al. (2023a); Deng et al. (2022); Hong et al. (2024b).
Based on the above analysis, we propose Decomposed Automation Correction (DAC), which corrects the generated SQL with the mistakes of entity linking and skeleton parsing as feedback. Schema linking denotes detecting the table and column names related to the question, where we use LLMs to directly generate relevant entities based on the user question and the database, following Li et al. (2023c). Skeleton parsing denotes generating the SQL skeleton corresponding to the user question, where we employ LLMs to generate the SQL skeleton according to the prior research Guo et al. (2024b; a). After generating the entities and skeleton, we take their inconsistencies with the initial SQL as feedback for the correction, as shown in Figure 1.
To validate DAC, we conduct experiments on three mainstream text-to-SQL datasets: Spider Yu et al. (2018), Bird Li et al. (2023b), and KaggleDBQA Lee et al. (2021). The experiments show that our method achieves an average improvement of $3.7\%$ compared to baselines, and a $1.3\%$ improvement over previous text-to-SQL automation correction methods, demonstrating the effectiveness of DAC.
@@ 2.2 Adaption to Text-to-SQL
We apply the discussion above to the text-to-SQL task, where we use entity linking and skeleton parsing as sub-tasks for automated correction. Entity linking refers to identifying relevant table and column names from the given database schema based on the user question Liu et al. (2021). Skeleton parsing denotes creating the corresponding SQL skeleton of the user question, which removes the table names, column names, and values in SQL Guo et al. (2024b). We use these two sub-tasks because previous work regards them as the most crucial capabilities for text-to-SQL Li et al. (2023a); Deng et al. (2022); Hong et al. (2024b).
@@ 3.1 SQL Generation
Before the correction, we first generate an initial SQL to be corrected. Following previous work (Xia et al., 2024), we use few-shot learning to generate SQL. We use BM25 to select demonstrations similar to the user question from the given demonstration pool, together with the database and the user question as the input. With the input, LLMs directly generate a single SQL as the initial result to be corrected.
@@ 3.2 Entity Linking
After obtaining the initial SQL, we aim to identify mistakes in the linked entities. To achieve this, it is essential to know the entities related to the user question, specifically the table and column names in the question-related database. Following previous work (Andrew et al., 2024; Goel et al., 2023), we use LLMs to select the question-related tables and columns. We employ the tokenized user question and the database as inputs and output the entity linking information in the format of List[Dict[str, str]] in Python, where the length of the list matches the number of tokens in the tokenized question. We use this format because the performance of entity linking generated in programmatic format by LLMs is superior to only generating entities (Li et al., 2023c).
@@ 3.3 Skeleton Parsing
After identifying the linked entities, we seek to parse the skeleton corresponding to the user question to identify mistakes in the initial SQL generation on the skeleton. Previous research indicates that LLMs can parse the SQL skeleton based solely on the user question since the skeleton corresponds to the syntactic logic of the question without requiring database information (Gu et al., 2023; Pourreza et al., 2024; Kothyari et al., 2023). Therefore, we use only the user question as the input and ask LLMs to parse an SQL skeleton corresponding to the question. Subsequently, we remove entities from the parsed SQL, since these entities could not correspond to the column names or table names in the original database, thus obtaining the SQL skeleton.
@@ 4.1 Settings
4.1.1 Datasets
We use Spider (Yu et al., 2018), Bird (Li et al., 2023b), and KaggleDBQA (Lee et al., 2021) to evaluate DAC. Spider is one of the most mainstream text-to-SQL datasets, which covers multi-table and multi-domain questions. Compared with Spider, Bird is more difficult, and databases are more complex and more near to the practical application. The difficulty of KaggleDBQA is close to that of Bird, but the scale of the database is larger [^2] We call KaggleDBQA as Kaggle for simplicity..
4.1.2 Metric
In this paper, following the previous works (Askari et al., 2024; Cen et al., 2024), we use execution accuracy (EX) to evaluate our method. EX refers to the proportion of data that predicted SQL execution results are the same as correct SQL execution results.
4.1.3 Models
We verify our methods on the instruction fine-tuning version of Llama3 (Meta, 2024), Deepseek-Coder (Deep.C.) (Guo et al., 2024c), and gpt-3.5-turbo-1106 (GPT-3.5) (Ye et al., 2023). Llama3 and Deepseek-Coder are the most mainstream general and code LLMs in the current open-source LLMs, respectively. While GPT-3.5 is one of the most popular closed-source LLMs. Therefore, our experimental LLMs can well cover different application scenarios.
4.1.4 Implementation Details
We employ few-shot inference for each step in DAC. Following the previous work (Chang & Fosler-Lussier, 2023), we use 5-shot for the inference. We use the Spider training set as the candidate demonstration pool and use BM25 to select the most similar demonstrations for each user question. For the entity linking step, we employ the labeled linking information of Spider collected by Liu et al. (2021), which is used to guide the generation format Madaan et al. (2023) and can be adapted to different databases as shown in Table 1. The prompts we used can be seen in the supplementary material.
@@ 4.2 Main Experiment
gpt-3.5-turbo Self-Debug (Chen et al., 2024) $71.1\rightarrow 72.2$-
SQL-CRAFT (Xia et al., 2024) $78.2\rightarrow 79.3$-
@@
gpt-4 Self-Debug (Chen et al., 2024) $73.2\rightarrow 73.6$-
EPI-SQL (Liu & Tan, 2024) $79.7\rightarrow 85.1$-
SQL-CRAFT (Xia et al., 2024) $79.7\rightarrow 85.4$ $53.3\rightarrow 55.2$
MAGIC (Askari et al., 2024) $78.6\rightarrow\bm{85.7}$-
@@ 4.4 Analysis
4.4.1 Is SQL with the correct entity and skeleton the correct SQL?
@@
4.4.2 Can LLMs effectively generate entity linking information?
@@
To validate the performance of LLMs in entity linking, we conduct experiments on the schema linking of the Spider dev set. The experimental results are shown in Table 5, from which we can see that: (i) In most metrics, the performance of entity linking based on LLMs is superior to that based on grams or fine-tuning methods Liu et al. (2021), demonstrating the effectiveness of LLMs in entity linking; (ii) Compared to code-specific LLMs, general LLMs exhibit better entity-linking performance in most metrics, indicating that general-purpose LLMs are more suitable for the entity-linking task due to the better instruction following.
4.4.3 Is the SQL skeleton generated not based on databases better than based on databases?
@@
4.4.4 What is the performance of DAC using oracle entities and skeletons?
@@
4.4.5 What is the performance of DAC on SQLs with different hardness?
@@ 5.1 Text-to-SQL
Research on the text-to-SQL task investigates how to generate corresponding SQL results based on given user questions and databases (Deng et al., 2022). The past research is mainly based on small-scale models Guo et al. (2019); Bogin et al. (2019); Wang et al. (2020); Scholak et al. (2021); Dou et al. (2022) like BERT Devlin et al. (2019) and BART Lewis et al. (2020). Currently, the methods based on LLMs have become the mainstream of the text-to-SQL task since their brilliant performance without fine-tuning (Hong et al., 2024b). One type of approach is to fine-tune LLMs to help them better adapt to the text-to-SQL task by directly using the text-to-SQL data (Zhang et al., 2024b; Li et al., 2024b; Pourreza & Rafiei, 2024), or merging the data of the tasks related to the text-to-SQL task (e.g., entity linking, knowledge generation) Sun et al. (2024); Hong et al. (2024a). Another type of approach uses LLMs by few-shot inference, achieving significantly better performance without the overhead of fine-tuning, like in-context learning Pourreza & Rafiei (2023); Gao et al. (2024); Chang & Fosler-Lussier (2023), self-consistency Ni et al. (2023); Zhong et al. (2023); Lee et al. (2024), and multi-agent Wang et al. (2024a); Xie et al. (2024).
@@ 5.2 Automated Correction
The automated correction approach first determines the correctness of the generated result after producing an answer, where if the answer is incorrect, it makes the necessary corrections (Pan et al., 2024). The most mainstream automated correction methods primarily focus on the code generation task, such as using feedback during training to correct errors (Ouyang et al., 2022; Feng et al., 2024; McAleese et al., 2024) or having the model judge and correct errors after inference (Chen et al., 2023; 2024; Wu et al., 2024; Zhang et al., 2024c). For the text-to-SQL task, automated correction methods pay more attention to the characteristics of the task itself, such as correcting based on execution results from the database (Li & Xie, 2024; Zhong et al., 2023; Xia et al., 2024) or generating SQL-related error messages (Liu & Tan, 2024; Askari et al., 2024).
However, previous work shows that LLMs struggle to directly identify SQL mistakes Zhang et al. (2024a); Kamoi et al. (2024). Therefore, we propose DAC to provide decomposed feedback from the sub-tasks of entity linking and skeleton parsing, thereby helping correct mistakes.
@@ Appendix A Prompts of DAC
The prompts of DAC is shown in Table 9, Table 10, Table 11, Table 12, and Table 13..